OpenStudy (luckycoins888):

FAN + MEDal! What is the volume of a cube, if the length of one side is 9 inches? A. 81 cubic inches B. 486 cubic inches C. 27 cubic inches D. 729 cubic inches

Directrix (directrix):

Volume = edge times edge times edge V = 9 ^3 V = 9 * 9 * 9 V = ? @luckycoins888

Directrix (directrix):

inches times inches times inches is cubic inches. How many cubic inches? Enter this in the Google search box and the Google Calculator will pop up with the product: 9 * 9 * 9 =